I need help with this problem:
An aircraft controller sees two aircraft flying at the same altitude on his screen. One, a Boeing 747 Jumbo, is headed Eastbound at 375 miles per hour. The other, a UFO, is located 57 miles directly south of the 747 and is directly northbound at 1750 miles per hour.
assuming they continue on this flight path how close (in miles) will the two aircraft come to each other?

Answer by rothauserc(4718)

First determine how much time it takes the UFO to cover the 57 miles
rate * time = distance
1750t = 57
t = 57/1750 = 0.03 hours
now let's see how far the jet flies in 0.03 hours
375* 0.03 = d
d = 11.25 miles
the UFO and jet come within 11.25 miles of each other